The Degree measure of an angle is 15 more than twice the measure of its compliment. Find the measure of the angle and its compliment.
Answer by Alan3354(69443)

The Degree measure of an angle is 15 more than twice the measure of its compliment. Find the measure of the angle and its compliment.
-----------------
The complement is 90 minus the angle (not compliment).
x-15 = 2(90-x)
x-15 = 180-2x
3x = 165
x = 55º